OpenStudy (anonymous):

Solve. (x + 2)^2 = 1

OpenStudy (anonymous):

x^2+4=1 x^2=5 x=2.5

OpenStudy (anonymous):

the answer choices are x=-3 and -1 x=3 and 1 x=-1 and 4 x=-4 and -3

OpenStudy (anonymous):

oh, wait, sorry, it's this:

OpenStudy (anonymous):

(x+2)^2-1=0 x^2+4x+4-1=0 x^2+4x+3=0 (x+1)(x+3)=0 x=-1 and -3

OpenStudy (anonymous):

do u get it?

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

s'ok :) do u get how i did it, though?

OpenStudy (anonymous):

yeah, I know how to do this, I just kept getting -3 and 1

OpenStudy (anonymous):

ohh, ok :)